How much water must be added to one gallon of 8% saline solution to get a 2% saline solution?
Answer by stanbon(75887)

How much water must be added to one gallon of 8% saline solution to get a 2% saline solution?
-----------------
Equation:
salt + salt = salt
0.08*1 + 0*x = 0.02(1+x)
---
Multiply thru by 100 to get:
8 + 0 = 2 + 2x
2x = 6
x = 3 gallons (amt. of water that must be added)
============================
Cheers,
Stan H.